OpenStudy (anonymous):

Help! Simplify: log(1/3) cubed root of 81 A. -4/3 B. 3/4 C. 4/3 D. 3

OpenStudy (kropot72):

Should the expression in the question be written as follows? \[\large \log_{3} \sqrt[3]{81}\]

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

Oh wait im sorry, its log(1/3) cubed root of 81

OpenStudy (kropot72):

The cube root of 81can be written as follows: \[\large \sqrt[3]{81}=3\times \sqrt[3]{3}\] Do you follow this step?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Would it be c?

OpenStudy (kropot72):

So now we can write the expression as follows: \[\large \log_{3} 3+\log_{3} 3^{\frac{1}{3}}=1+\frac{1}{3}\]

OpenStudy (anonymous):

okay I follow

OpenStudy (kropot72):

Good! So is choice c correct?

OpenStudy (anonymous):

Yes

OpenStudy (kropot72):

Your choice is correct.

OpenStudy (anonymous):

Thank you!

OpenStudy (kropot72):

You're welcome :)